A certain university produced 895 licensed teachers after they successfully passed the Licensure Examination for Teachers (LET). The professional Regulation Commission (PRC) records show that the examinees form that university posted an average rating of 85 with standard deviation of 2. If 36 successful examinees are selected at random, what is the probability that their ratings fall between 84 and 86?


1
Expert's answer
2022-03-17T06:10:03-0400

Solution

Population mean "\mu =85"

Population s.d "\sigma=2"


(a) Probability that ratings fall between 84 and 86


"Z=\dfrac{X-\mu}{\sigma}"


"X_1=84 ~~and ~~X_2=86"


"Z_1=\dfrac{84-85}{2}=-0.5"


From normal distribution tables

"P(Z of -0.5) =0.30854"


"Z_2=\dfrac{86-85}{2}=0.5"


From normal distribution tables

"P(Z of 0.5)=0.69146"

For values above "86"

"P=1-0.69146=0.30854"


Probability that the values falls between 84 and 86


"=1-(0.30854+0.30854)"


"=0.38292"